OpenStudy (frankypoo):

Find the derivative: x√(100-x^2) please help.

OpenStudy (frankypoo):

\[x \sqrt{100-x ^{2}}\]

OpenStudy (anonymous):

use product rule and get \[\sqrt{100-x^2} + x^2/\sqrt{100-x^2}\]

OpenStudy (anonymous):